OpenStudy (anonymous):

Agebra help please? f(3) when f(x)=2x+2 y+10=2y+1

OpenStudy (anonymous):

If f(x)= f(a), what is x?

OpenStudy (anonymous):

I know that one

OpenStudy (anonymous):

Hold on I'm trying to post a third equation

OpenStudy (anonymous):

Those two are 8 and 9

OpenStudy (anonymous):

its not working..lol

OpenStudy (anonymous):

ill try to write it out

OpenStudy (anonymous):

OK, WILL WAIT

OpenStudy (anonymous):

f^-1 (4) (3x-4)/5

OpenStudy (anonymous):

im trying to write an equation but the button isnt working

OpenStudy (anonymous):

ok. lets sort the first one first. we will deal with thereafter. u will write it in words. f(x) =f(a), what do u think is x?

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

Wouldn't it be f(3)2x+2 = 6x+2 = 8x?

OpenStudy (anonymous):

ok, x=a, f....(a) and then f...(x) , so x = a. what this means is in place of x, put a, now in your first question, a =3, so in place of x , what do u put?

OpenStudy (anonymous):

I put a

OpenStudy (anonymous):

u must put 3 in place of x,

OpenStudy (anonymous):

because its f(3)

OpenStudy (anonymous):

I did...that's how I got 8x

OpenStudy (anonymous):

f(3)2x+2 = 6x+2 = 8x

OpenStudy (anonymous):

I have to find which 2 problems are alike..

OpenStudy (anonymous):

ok, if we put 3 IN PLACE OF x, it means NO MORE x, x is out, now there is 3.

OpenStudy (anonymous):

i meant 8a

OpenStudy (anonymous):

doesnt matter

OpenStudy (anonymous):

we have 8..next problem?

OpenStudy (anonymous):

So it wil bel 2(3) +2

OpenStudy (anonymous):

next pproblem , put LIKE TERMS together,terms with y on their own side

OpenStudy (anonymous):

its 9

OpenStudy (anonymous):

i need the other one though

OpenStudy (anonymous):

I told you I had these 2 solved already

OpenStudy (anonymous):

--> f^-1 (4) (3x-4)/5 <--

OpenStudy (anonymous):

YEP. 2y-y=10-1, y=9

OpenStudy (anonymous):

ok, type it in words

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

--> f^-1 (4) (3x-4)/5 <--

OpenStudy (anonymous):

--> f^-1 (4) (3x-4)/5 <--

OpenStudy (anonymous):

--> f^-1 (4) (3x-4)/5 <--

OpenStudy (anonymous):

^^^

OpenStudy (anonymous):

its weird..the f has a small -1 above it

OpenStudy (anonymous):

yes, a power of minus 1 means 1 over......so it means the inverse of the function f. it means 1 divided by that equation

OpenStudy (anonymous):

1 over that whole equation, unfortunately the button is not working so i can't tell head from tail.

OpenStudy (anonymous):

is it working for you?

OpenStudy (anonymous):

The problem is f^-1 (4) (3x-4)/5 3x-4 is the numerator and 5 is the denominator of the fraction. 4 is the function and -1 is the exponent of f

OpenStudy (anonymous):

does that help at all?

OpenStudy (anonymous):

oh, ok. oh, i see. ha ha ha. u have a funny way of writing when the button z not working! ok, in place of x, we put 4, so that we will have for the NUMERATOR: (3(4)-4)= 12-4=8 DENOMINATOR IS ALWAY 5, so we will have f^(-1) 8/5. this means the inverse of 8/5. if the inverse of 4/3 = 3/4 what is the inverse of 8/5?

OpenStudy (anonymous):

@bbbbbbbbbbbbbbbbbb ?

OpenStudy (anonymous):

sorry im back

OpenStudy (anonymous):

How is it 5/8 though? I have to find out which 2 problems are the same and if its 5/8 then none of them are the same

OpenStudy (anonymous):

MAY BE IF U CAN TAKE A PIC OF THE QUESTION AND UPLOAD.

OpenStudy (anonymous):

idk how to take screenshots

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

wc windows r u using

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

uh

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

how can i tell?

OpenStudy (anonymous):

ok, use your 4n

OpenStudy (anonymous):

whats 4n?

OpenStudy (anonymous):

a phone with a camera

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

good idea!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

will w8

OpenStudy (anonymous):

Number 2/II

OpenStudy (anonymous):

good gal/boy? the question reads wc of the following what?

OpenStudy (anonymous):

Which of the following statements have the same result? Explain each step in solving each one.

OpenStudy (anonymous):

ya @bbbbbbbbbbbbbbbbbb . we have done everything alright. the answr ther is as i said, 5/8. i showed u the steps, so, none of them brings same result .

OpenStudy (anonymous):

Okay, thank you for your help :)